OpenStudy (anonymous):

Trig identity: 1/sin^2X + 1/cos^2X = 1/sin^2X cos^2X

OpenStudy (anonymous):

Are you being asked to PROVE or SHOW that this identity is true?

OpenStudy (anonymous):

\[1divsin^2x+1divcos^2x=(\cos^2x+\sin^2x)divsin^2xcos^2x\] but cos squared plus sin squared equals one QED